LUCIFER Technical Observing, prep for Partner Science:
UT 2010 Jun 06

Observer: DThompson
Telescope Operator: SAllanson
Instrument Support: JMorris
Telescope Support: JLittle
SW support: JKraus on call

Summary

  • Aside from MOS unit being locked out, telescope and instrument worked well
  • LUCIFER computer found to be 30s ahead of TCS computers (that all agree with each other).
  • Twilight flats in BrG, H2, FeII, Pbeta, and Ks obtained
  • LBCR i-band flats obtained for yesterday's data
  • Gathered NB illumination data in BrG
  • Verified offsets work as expected
  • Verified collimation model is OK
  • Made new pointing model: LUCIFERSX.20100606.ptmod, set as default (4.1s rms)
  • Collected field aberration data for Doug/Andrew
  • Obtained Ks, H2, and BrG images of Messier 57.

Preparations

Mask exchange did not take place due to excessive temperatures in the motor control electronics. Peter has shut out MOS movements to protect the instrument. Imaging only.

Details

Files: luci.20100606.NNNN.fits

Twilight flats:
  • N3.75+mirror, no mask in use, ROT@341 (unless otherwise noted).
  • 01:53 0002-0011 BrG+clear 2.0s 11kDN
  • 01:58 0012-0021 H2+clear 2.0s 12kDN
  • 02:02 0022-0031 BrG+clear 2.0s 9.5kDN
  • 02:04 0032-0041 BrG+clear 2.0s 9.4kDN ROT@161 (check changes in gradients)
  • 02:16 0042-0051 BrG+clear 2.0s 6.9kDN
  • 02:18 0052-0061 BrG+clear 2.0s 6.3kDN ROT@161
  • 02:21 0062-0071 H2+clear 2.0s 6.7kDN
  • 02:22 0072-0081 FeII+clear 2.0s 12kDN
  • 02:24 0082-0091 Pbeta+clear 2.0s 14kDN
  • 02:29 0092-0101 BrG+clear 2.0s 2.7kDN
  • 02:31 0102-0111 H2+clear 2.0s 2.4kDN
  • 02:33 0112-0121 FeII+clear 2.0s 4.4kDN
  • 02:35 0122-0131 Pbeta+clear 2.0s 4.8kDN
  • 02:38 0132-0141 clear+Ks 2.0s 8kDN
  • 02:44 0142-0151 clear+Ks 2.0s 3.5kDN

Correct initial collimation/pointing:
  • Initial IE=-47.25 CA=0.00 IA=-400
  • Img 152: IE=-59 CA=-24
  • Collimate on BS9149 on axis, EL~57 (sky too bright still)
  • 03:01 Letting LBCR take some i-band flats

LBC Red I-BESSEL Sky Flats
  • 03:03 (30 minutes after sunset) counts are ~10,000 in 1 sec (chip 2 only!)
  • 03:06 set of 5 3 sec flats. 030635 - 030840 (some stars but dithered, ~14k)

Check initial collimation/pointing:
  • Initial IE=-47.25 CA=0.00 IA=-400 (reloaded when re-authorized?) -0.5mm M1 focus
  • Img 153: IE=-65 CA=-27
  • Collimate on BS9149 on axis, EL~57 Seeing ~0.5"

Gather NB illumination correction data
  • 0154 test exposure for scaling
  • 03:34 Execute D25x_illum1.scr on BS9149 with off-axis guide star, EL~55. 6x10.0s int, BrG. Files 0155-0179
  • 04:13 Execute D25x_illum2.scr on BS9149 with off-axis guide star, EL~51. 6x10.0s int, BrG. Files 0180-0204
  • Guider looked decent (~0.5-0.7 arcsec FWHM) on off-axis star throughout, on-axis images clearly elongated consistently from upper-left to lower-right.

Run science-type dice5 pattern to check that things still work
  • Check repeated positioning - Illum scripts repeat center position and flexure compensation was OFF on the UI and not started in the scripts.
  • Visually confirmed the ABS/DETXY offsets were executed as expected in the illum scripts.

Note in script-processing window if DETXY offsets are noted as RA and DEC
Confirmed that the script is set up for DETXY offsets and the xterm where the script was executed reports this, but the offsets are listed as dRA and dDE:
Executing observation element 5 of 25
Using absolute offsets in DETXY
offsetting the telescope by -71.73'' RA and -36.2'' DEC.
No rotator angle offset set for this position. Using :0.0
offsetting the telescope finished
beginning read...
but note that the offsets listed in the script are: OFFSET =  -71.73  -36.20 so the offsets are incorrectly reported in the xterm. Marcus emailed.

Check existing collimation model
  • Collimate on BS9155, EL~57 (0205, 0206). Sources definitely elongated in 0205 but good in 0206.
  • Check low EL on BS9138 in track mode (0207, blind filter in), EL~13 M1 hit Z and Y limits (parallel transfer to back off)
  • 05:07 Collimate again on BS9155, img 0208
  • 05:12 BS9143 at EL~24 (0209) looks good
  • 05:15 BS9166 at EL~69 (0210) looks OK, elongated but compact
  • ptcorr here, file 0211 IE now -75, CA now -38, then collimate

Get data for new pointing model
  • /home/telescope/20100606.ptlog started 05:38, done 06:38

01 ACT0356 0214 02 ACT0405 0215 03 ACT0399 0216 04 ACT0339 0217
05 ACT0290 0218 06 ACT0188 0219 unwrap AZ 07 ACT0245 0220 08 WT10_090 0221
09 WT10_160 0222 10 ACT0266 0223 11 WT10_217 0224 12 WT10_319 0225
13 ACT0332 0226 14 ACT0287 0227 15 WT10_270 0228 16 skipped, no star
17 ACT0194 0229 18 WT10_092 0230 19 WT10_139 0231 20 WT10_281 0232
21 WT10_243 0233 22 WT10_222 0234 23 WT10_292 0235 24 ACT0328 0238 came back to do this one
25 ACT0401 0236 26 WT10_407 0237 27 ACT0411 0239 28 ACT0343 0240

  • New LBT Pointing model, 2010 Jun 6 LUCIFERSX.20100606.ptmod (set as default) RMS=4.1 arcsec #
  IA       -398.1615    10.26739
  IE        -84.5313     1.49660
  NPAE      +21.7646    11.91240
  CA        -61.8785    14.62254
  AW         -0.5047     0.92274
  AN        +21.7423     0.92739
  TX         +8.3799     0.97103

Collect field aberration dataset, StoneM field
  • in /OldRepository/AGW_Data/20100606
  • copied /home/LBTO/iraf/!field_StoneM.cl to this directory
  • pointto StoneMagw gs=0 Tel="ACTIVE" PA=0, allow to collimate
  • cl < field_StoneM.cl started at 06:57 but Doug called 1min after starting to say it is not the right script.
  • use field_BS9177.cl instead, modified for LFBG by Doug
  • 07:23 started field_BS9177.cl (local copy) after correcting pointing and collimating here.
  • 08:59 script finished, 01:46 to execute.

GStar wfscimage#
0 7-10
40 11-13
0 14-16
57 17-19
0 20-22
46 23-25
0 26-28
13 29-31
0 32-34
30 35-37
0 38-40
54 41-43
0 44-46
32 47-49
0 50-52
14 53-55
0 56-58
41 59-61
0 62-64
57 65-67
0 68-70
46 71-73
0 74-76
13 77-79
0 80-82
30 83-85
0 86-88
50 89-91
0 92-94

Image: Messier 57
  • 9x10.0s Ks using Ks_ Files 0262-280 (saved in normal mode)
  • 9x10.0s Ks Files 0281-298
  • 299 30s H2 for scaling (use 2x60s)
  • 2x60.0s H2 Files 300-317
  • 2x60.0s BrG Files 318-326, collimation obviously suffering from twilight

Repeat measurement of focal plane scale, orientation
  • Use M57, data spans EL 87 to 61

Darks for these data:
  • 10x2.0s Files 0327-0331
  • 9x10.0s Files 0332-0336
  • 2x60.0s Files 0337-0341

Not Done:
  • Through-slit images and FF spectra of all masks to compare German-cut to URIC-cut masks. MOS locked out
  • Re-test updatePointingReference functionality if re-implemented in LUCIFER SW. Functionality not re-implemented yet
  • June 2010: Test GCS guider camera 2x2 binning mode. Torsten not reachable by phone.

-- DavidThompson - 05 Jun 2010
Topic revision: r10 - 09 Jul 2010, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback